OpenStudy (anonymous):

How do you work the intercepts out in solving the following equation graphically y=x3(cubed)-7x2(squared) - 14x +120

OpenStudy (anonymous):

to find y intercept set x= 0 to find x intercept set y = 0 and solve for x

OpenStudy (anonymous):

ok to find the intercepst you have to make y=0 and solve for x to find x intercepts and x=0 and solve for to find y intercepts.

OpenStudy (anonymous):

can you do the example and show me workings please

OpenStudy (shubhamsrg):

x=0 gives y intercept => 120 y=0 gives x intercept x^3 - 7x^2 -14x +120 =0 x^3 + 4x^2 - 11x^2 - 44x + 30x + 120=0 => (x^3 - 11x^2 +30x) + (4x^2 -44x + 120)=0 can you do now?

OpenStudy (shubhamsrg):

or you may use rational root theorem..

OpenStudy (anonymous):

Sorry shubhamsrg i dont get the second line. how did you end up with 4x^2 - 11x^2 etc

OpenStudy (anonymous):

Integer factor of the coefficient of the first term and integer factor of the last term which is the constant*

OpenStudy (anonymous):

rational root theorem all possible roots of the equation = ±p/q

OpenStudy (anonymous):

x^3 - 7x^2 - 14x + 120 ok to find y intercept set all values of x to 0 so y intercept = 0 - 0 - 0 + 120 now to find your x intercept set y = 0 and solve for x

OpenStudy (anonymous):

Y is easy, need someone to show me how to do x. Need some workings

OpenStudy (anonymous):

I know the answer is (-4,0) (5,0) and (6,0)

OpenStudy (anonymous):

well what did you learn when you have a 3 degree polynomial to solve for x when y = 0 ? There are a couple of ways you can do it 1) draw the graph and see where x is 0 2) Use the quadratic equation and plug in the values The easiest is by using the quadratic formula